# Towards a Modern and Lightweight Rendering Engine for Dynamic Robotic Simulations

### v1

@@ V Experimental results
To evaluate the computational performance of the complete simulation framework, AMBF and AMBF-Vulkan were run on an Ubuntu 22.04 workstation with an NVIDIA GeForce RTX 4090 GPU. Two Blender scenes were used to evaluate the frameworks. Scene 1 includes two sets of da Vinci Patient Side Manipulators (PSMs) and a training suturing phantom taken from [2, 9], and has a total of 223K triangles. Scene 2 includes a model of a full da Vinci Si that was acquired from DistroSquid [^1] https://www.turbosquid.com and has a total of 550K triangles. Each scene was exported from a Blender file into ADF and glTF formats and then loaded to AMBF and AMBF-Vulkan, respectively. For every test, all the geometry was kept within the clipping plane, and no culling operations were applied to reduce the triangle count.
V-A AMBF-Vulkan computational performance
Figure 7 shows the GPU processing time for each stage in the rendering pipeline of AMBF-Vulkan as the total number of triangles increases. To easily augment the number of triangles, we duplicated the objects in Scene 1 so that the number of triangles would grow exponentially at every test. Using NVIDIA Nsight Graphics GPU Trace, five frame captures of each scene were profiled. The computational times for the following tasks were recorded: TLAS building, the main render pass, the post-processing pass, and the ImGui pass.
Figure 7: GPU computation time measurements of segments of Vulkan’s render loop. Measurements were taken from scenes comprised of increasingly greater triangle counts.
@@ V-A AMBF-Vulkan computational performance
V-B Quantitative comparisons between AMBF-Vulkan and AMBF
Engine AMBF AMBF-Vulkan
@@ V-B Quantitative comparisons between AMBF-Vulkan and AMBF
TABLE I: Mean GPU usage on the render loop’s main pass in AMBF and AMBF-Vulkan measured in ms.
Table I shows the main pass GPU computation time of both AMBF-Vulkan and AMBF while rendering scene 1 and scene 2. Similarly to the previous test, NVIDIA Nsight Graphics GPU Trace was used to capture twenty frames of each scene for each engine. For each frame capture, we measured the computational time for the main render pass, which only consists of the drawing and shading of the scene’s geometry.
The results show that AMBF-Vulkan has an average speed increase of 6.1 and 8.5 times over AMBF for scene 1 and scene 2, respectively. Additionally, AMBF-Vulkan’s measurements showed less variance. AMBF’s computation times increased from scene 1 to scene 2, which is expected as scene 2 contains more triangles and more textures to render. On the other hand, AMBF-Vulkan used roughly the same computation time for both scenes. This is likely due to the GPU computation being so fast that the application is CPU-bound.
V-C Qualitative comparisons between AMBF-Vulkan and AMBF
Figure 8: Scene 2 rendered from the same viewpoint in AMBF (top) and AMBF-Vulkan (bottom) for qualitative comparison.
Scene 2 was rendered by both engines from the same viewpoint with two lights and shadows enabled to perform qualitative comparisons (see Fig. 8). Improvements made to the reflection model in AMBF-Vulkan enable an increased visual clarity of the model’s details and an overall more physically realistic image. Additionally, the ray-traced shadows in AMBF-Vulkan are more precise than the AMBF shadows generated with shadow maps. Lastly, the jagged anti-aliasing artifacts at the edges of the geometry in AMBF are virtually non-existent in AMBF-Vulkan.
@@ VI Discussion and Future Work
This paper presents the initial infrastructure required to modernize the AMBF simulator with a modern graphics API capable of taking advantage of GPU hardware advancements. Most of the features presented in this paper serve as groundwork for future iterations of the engine. For example, vertex buffer addressing, a PBR reflection model, and the acceleration structure build system are all key preliminaries to a full ray-tracing pipeline. Additionally, incorporating a post-processing render pass will easily enable us to implement more advanced image processing effects, such as depth-of-field and bloom. Lastly, integration with the Dear ImGui library is the first step toward a custom scene editor that can enable modifying scenes at runtime.
The planned development of the new engine is focused on three major goals. (1) The full migration of existing AMBF capabilities to create a more performant, unified framework for robotics simulation. (2) Further development of the engine’s ray-tracing pipeline to produce physically based and photorealistic simulations. (3) Improvement of the framework’s user experience and tooling to minimize the barrier to entry and increase simulation development productivity.
Other potential areas of research and development include GPU-accelerated soft-body dynamics, volumetric smoke simulation from cutting tools, fluid simulation for pooling bodily fluids, physically based endoscope lenses, reduced visual clarity from lens condensation, sub-surface scattering material models for realistic tissue rendering, and image-based lighting from 360°  abdomen-interior high-dynamic-range images. Overall, the proposed Vulkan-based simulation engine lays the foundation to produce more performant and physically accurate simulations in multiple research domains.
@@
For more information, visit the project repository at https://github.com/AMBF-Vulkan-repositories/AMBF-Vulkan

### v2

@@ V Experimental results
To evaluate the computational performance of the complete simulation framework, AMBF and FIRE-3DV were run on an Ubuntu 22.04 workstation with an NVIDIA GeForce RTX 4090 GPU. Two Blender scenes were used to evaluate the frameworks. Scene 1 includes two sets of da Vinci Patient Side Manipulators (PSMs) and a training suturing phantom taken from [3, 10], and has a total of 223K triangles. Scene 2 includes a model of a full da Vinci Si that was acquired from DistroSquid [^1] https://www.turbosquid.com and has a total of 550K triangles. Each scene was exported from a Blender file into ADF and glTF formats and then loaded to AMBF and FIRE-3DV, respectively. For every test, all the geometry was kept within the clipping plane, and no culling operations were applied to reduce the triangle count.
V-A Computational performance of FIRE-3DV
Figure 7 shows the GPU processing time for each stage in the rendering pipeline of FIRE-3DV as the total number of triangles increases. To easily augment the number of triangles, we duplicated the objects in Scene 1 so that the number of triangles would grow exponentially at every test. Using NVIDIA Nsight Graphics GPU Trace, five frame captures of each scene were profiled. The computational times for the following tasks were recorded: TLAS building, the main render pass, the post-processing pass, and the ImGui pass.
Figure 7: GPU computation time measurements of segments of FIRE-3DV’s render pipeline. Measurements were taken from scenes comprised of increasingly greater triangle counts.
@@ V-A Computational performance of FIRE-3DV
V-B Quantitative comparisons between AMBF and FIRE-3DV
Engine AMBF FIRE-3DV
@@ V-B Quantitative comparisons between AMBF and FIRE-3DV
TABLE I: Mean GPU usage on the render loop’s main pass in AMBF and FIRE-3DV measured in ms.
Table I shows the main pass GPU computation time of both FIRE-3DV and AMBF while rendering scene 1 and scene 2. Similarly to the previous test, NVIDIA Nsight Graphics GPU Trace was used to capture twenty frames of each scene for each engine. For each frame capture, we measured the computational time for the main render pass, which only consists of the drawing and shading of the scene’s geometry.
The results show that FIRE-3DV has an average speed increase of 6.1 and 8.5 times over AMBF for scene 1 and scene 2, respectively. Additionally, FIRE-3DV’s measurements showed less variance. AMBF’s computation times increased from scene 1 to scene 2, which is expected as scene 2 contains more triangles and more textures to render. On the other hand, FIRE-3DV used roughly the same computation time for both scenes. This is likely due to the GPU computation being so fast that the application is CPU-bound.
V-C Qualitative comparisons between AMBF and FIRE-3DV
Figure 8: Scene 2 rendered from the same viewpoint in AMBF (top) and FIRE-3DV (bottom) for qualitative comparison.
Scene 2 was rendered by both engines from the same viewpoint with two lights and shadows enabled to perform qualitative comparisons (see Fig. 8). Improvements made to the reflection model in FIRE-3DV enable an increased visual clarity of the model’s details and an overall more physically realistic image. Additionally, the ray-traced shadows in FIRE-3DV are more precise than the AMBF shadows generated with shadow maps. Lastly, the jagged anti-aliasing artifacts at the edges of the geometry in AMBF are virtually non-existent in FIRE-3DV.
@@ VI Discussion and Future Work
This paper presents the initial infrastructure required to modernize a dynamic simulation engine, such as AMBF, with a modern graphics API capable of taking advantage of GPU hardware advancements. Most of the features presented in this paper serve as groundwork for future iterations of the engine. For example, vertex buffer addressing, a PBR reflection model, and the acceleration structure build system are all key preliminaries to a full ray-tracing pipeline. Additionally, incorporating a post-processing render pass will easily enable us to implement more advanced image processing effects, such as depth-of-field and bloom. Lastly, integration with the Dear ImGui library is the first step toward a custom scene editor that can enable modifying scenes at runtime.
The planned development of FIRE-3DV is focused on three major goals: (1) the full migration of existing AMBF capabilities to create a more performant, unified framework for robotics simulation, (2) further development of the engine’s ray-tracing pipeline to produce physically based and photorealistic simulations, and (3) improvement of the framework’s user experience and tooling to minimize the barrier to entry and increase simulation development productivity.
Other potential areas of research and development include GPU-accelerated soft-body dynamics, volumetric smoke simulation from cutting tools, fluid simulation for pooling bodily fluids, physically based endoscope lenses, reduced visual clarity from lens condensation, sub-surface scattering material models for realistic tissue rendering, and image-based lighting from 360°  abdominal high-dynamic-range images. Overall, the proposed Vulkan-based simulation engine lays the foundation to produce more performant and physically accurate simulations in multiple research domains.
Acknowledgments and Disclosures
This work was supported in part by NSF AccelNet awards OISE-1927354 and OISE-1927275.
@@
For more information, visit the project repository at
https://github.com/FIRE-3DV-repositories/FIRE-3DV.